OpenStudy (anonymous):

how do i write an equation if im giving... through (10,2) and (2,-2)

OpenStudy (saifoo.khan):

use the formula.

OpenStudy (anonymous):

do u have the formula

OpenStudy (chaise):

m = y-y/x-x m = -2-2/2-10 m = 0.5 y=mx+b 2=0.5x10+b 2=5+b -3=b b=-3 y=mx+b y=0.5x-3